OpenStudy (chriss):

\[=\int\limits_{-1}^{1}f(x)g(x)dx\] \[f(x)=x\] \[g(x)=x ^{2}-x+2\] find a. b. ||f|| c. ||g|| d. d(f,g)

OpenStudy (zarkon):

where are you stuck?

OpenStudy (chriss):

Actually, I think I'm on a line it solving it now. I'll put my results in in a few minutes if you wouldn't mind verifying it for me.

OpenStudy (zarkon):

ok

OpenStudy (chriss):

Here's my a and b... if I have those right, I assume I've got this figured out... \[a. \int\limits_{-1}^{1}-x ^{3}+x ^{2}-2x\] \[b. \frac{2}{3}\]

OpenStudy (zarkon):

I don't get that

OpenStudy (zarkon):

\(<f,g>=-\frac{2}{3}\)

OpenStudy (zarkon):

\[\|f\|=\frac{\sqrt{6}}{3}\]

OpenStudy (zarkon):

\[<f,g>=\int\limits_{-1}^{1}x\times(x^2-x+2)dx\]

OpenStudy (chriss):

so for a. <f,g> since it's just defined as integral(f(x)(g(x))dx don't I just multiply my two formulas to get my integrand?

OpenStudy (zarkon):

\[\|f\|=\sqrt{<f,f>}=\sqrt{\int\limits_{-1}^{1}x\times x\,dx}\]

OpenStudy (chriss):

except when I integrated that, I got 2/3 not -2/3

OpenStudy (zarkon):

\[<f,g>=\int\limits_{-1}^{1}x\times(x^2-x+2)dx=\int\limits_{-1}^{1}(x^3-x^2+2x)dx\]

OpenStudy (chriss):

oh... I see had a typo... f(x) is -x not x... my bad

OpenStudy (zarkon):

oh...ok

OpenStudy (chriss):

so I get 2/3 for a, now with the typo corrected, is my answer for b correct?

OpenStudy (zarkon):

no...my answer is still valid since -x*-x=x*x=x^2

OpenStudy (chriss):

yeah, I gotta recalculate b... I left out entire steps that I see now... lemme see if I can get what you got.

OpenStudy (zarkon):

\[\|f\|=\sqrt{<f,f>}=\sqrt{\int\limits_{-1}^{1}(-x)\times (-x)\,dx}\] \[\sqrt{\int\limits_{-1}^{1}x^2\,dx}=\sqrt{\left.\frac{x^3}{3}\right|_{-1}^1}=\sqrt{\frac{1}{3}+\frac{1}{3}}=\sqrt{\frac{2}{3}}=\frac{\sqrt{6}}{3}\]

OpenStudy (chriss):

ok... that's what I get now too. thanks!

OpenStudy (zarkon):

good

OpenStudy (zarkon):

I need to go...but here is what I got for c and d c)\[\frac{4\sqrt{165}}{15}\] d)\[\frac{\sqrt{2490}}{15}\]

OpenStudy (chriss):

great. thanks again!